Can you please explain the differencs between argument & principal value of a complex number?

Grade:12

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear nilandri sarkar

principle value of an argumnet is a perticular value of argumnet that lies between  

 -pi< prinipal argu <= +pi

 for example

 let a complex number  z=1+i

                                    =√2(cos pi/4 +i sin pi/4)

                                     = √2(cos (2npi+pi/4) +i sin (2n pi pi/4))   here n is intiger

             here argument can be any value  for differen value of n

             but principle argument value  is pi/4      lies -pi< prinipal argu <= +pi

 

Please feel free to post as many doubts on our discussion forum as you can. If you find any question Difficult to understand - post it here and we will get you the answer and detailed solution very quickly.
 We are all IITians and here to help you in your IIT JEE preparation.

 All the best.
 
Regards,
Askiitians Experts
Badiuddin